OpenStudy (anonymous):

using quadratic formula... The speed of the current in a canal is 4mph. It takes twice as long for a houseboat to go 18 miles upstream as it does for it to go downstream. Find the speed of the boat in still water.

OpenStudy (anonymous):

\[\frac{ 36\pm \sqrt{(-36)^{2}-4(8)(-128)} }{ 16 } \]

OpenStudy (anonymous):

\[\frac{ 36\pm \sqrt{(-36)^{2}+4096} }{ 16 }\]

OpenStudy (anonymous):

\[\frac{ 36\pm \sqrt{1296+4096} }{ 16 }\]

OpenStudy (anonymous):

\[\frac{ 36\pm 73.1 }{ 16 }\]

OpenStudy (anonymous):

\[6.8 or -2.3\]

OpenStudy (anonymous):

44 8.5.6

OpenStudy (anonymous):

\[(15.8+2x)(25.5+2x)=385.3\]

OpenStudy (anonymous):

hi